OpenStudy (anonymous):

solve the system by the elimination method. 5x+3y=-11 7x-2y=17

OpenStudy (lalaly):

multiply the first by 2 and the sec by 3 you get 10x+6y=-22 21x-6y=51 add these two equaitons you get 31x=29 so x=29/31 substitute in any one of the equations to find y

OpenStudy (anonymous):

is this an ordered pair

jhonyy9 (jhonyy9):

y=(-11 -5(29/31))/3 =(-11 -145/31)/3 = (-341-145)/93 = - 486/93

jhonyy9 (jhonyy9):

than x = 29/31